OpenStudy (ujjwal):

If y=x^y then dy/dx= a)\[\frac{y^2}{x(1+ln y)}\] b)\[\frac{y^2}{x(1-lny)}\] c)\[\frac{2x}{y(1+ln x)}\] d)\[\frac{2y}{x(1+lny)}\]

OpenStudy (amistre64):

can we do implicit? or cange over to logs?

OpenStudy (ujjwal):

we can do anything we like

OpenStudy (amistre64):

y = x^y log(y) = log(x^y) log(y) = y log(x) 1/y y' = y' log(x) + y/x

OpenStudy (anonymous):